OpenStudy (anonymous):

Solve the equation for y in terms if x, and replace y with function notation f(x). Then find f(2) y+5x^2=5 f(x)= f(2)=

OpenStudy (bahrom7893):

f(x)=-5x^2+5

OpenStudy (bahrom7893):

f(2)=-5*2^2+5=-5*4+5=-15